Two unlike parallel forces of magnitude 400n and 150n are acting in such a way that their lines of action are 160mm apart. Determine the magnitude of it's resultant force and the point at which it act


1
Expert's answer
2021-03-12T09:29:18-0500
"F=400-150=250\ N\\ 400x=150(160-x)\\x=43.6\ mm"